wiki:UC20200602

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 06/02/2020

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), John Bally (CU), Moire Prescott (NMSU), Mukremin Kilic (OU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Sarah Tuttle (UW), Ben Williams (UW), Misty Bentz (GSU), Kal Kadlec (UW), Jamey Eriksen (APO)


We started by introducing Jamey Eriksen, the new Director of Site Operations, who will begin officially on June 8 (Mark Klaene will retire on August 1). Jamey earned a Masters degree in Astronomy from San Diego State University working with Bob Leach. He has since worked at several different observatories (CFHT, Palomar) and in industry (e.g. Ball Aerospace), and comes to us from Palomar Observatory. Welcome Jamey!

User feedback and comments from institutional representatives:

  • Oklahoma - nothing to report
  • NMSU - nothing to report
  • Colorado - nothing to report
  • Wyoming - nothing to report
  • Seattle - nothing to report
  • Washington - nothing to report
  • NAPG - no report
  • JHU - nothing to report, but eager to hear about the plans for Q3
  • UVa - nothing to report, but user for tonight wants to know whom to call for the password (answer: call the site)
  • Georgia State - nothing to report
  • FGCU - no report

Mark asked whether anyone envisions any problems with users utilizing their time? We confirmed that everyone can observe from home.


Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 4/29/2020 - 5/27/2020

0) Overview

APO has been closed for all observations due to NM Government restrictions. There has been no observing and absolute minimal staff on site for absolute minimal time. Non essential systems were powered down or shut off, the telescopes were safed, and instruments were kept cold and operating, with the exception of NICFPS (see below).

With a change in the state restrictions in late May and with a comprehensive reopening plan in hand, we submitted our plan to NMSU for approval to reopen. Having obtained approval of this plan, the APO executive committee decided to reopen APO starting on the night of June 1 for observations. Fortunately, we were able to start preparing for those activities early and that is ongoing; hopefully this will lead to a smooth return to night operations.

Please note that we are required to operate at no more than 25% staffing level. We have decided to restrict all visitors to the site at this time and restrict travel. We also remain under travel restrictions from the State of New Mexico and NMSU at this time.

1) Telescope

The telescope has been returned to operational status and readied for June 1 observations. Engineering tasks will be the first order of business to verify telescope performance.

2) Instruments

DIS is powered on and has been kept cold via cryotiger chillers; it is operational. The scattered light in the blue camera is high, rendering it generally unusable. We should have an update on the scattered light in the red camera by the time of the Users Committee meeting.

TripleSpec is powered on, cold and operational.

Agile is powered on and ready. We are currently reverted back one software version to correct connectivity issues.

Echelle is powered on, cold and operational. The red IOL is holding; the blue IOL got worse last month (possibly due to warming weather).

NIC-FPS is warm and powered off. It will be cooled after we can open it up and replace the ion pump.

ARCTIC is powered on, kept cold via a cryotiger chiller, and operational.

3) ARCSAT

ARCSAT is being brought back on line and should be ready for on-sky observations on June 1.


Additional telescope and instrument discussion:

Tonight is our first night of science observations; we did some on-sky engineering last night after the preparatory work of late last week to get the telescope back up and running. ARCSAT was on sky over the weekend for engineering and is now ready for science. There are extensive protocols in the workplace for maintaining staff health and safety. The main thing to stress is that we are operating at a 25% staffing level, which means we are shuffling which people can go in during the day. That, plus the wearing of masks and social distancing, may limit our ability to get some things accomplished. Responses to emails and addressing some problems will be slower. UC reps should ask their users to be extra patient with us.

The DIS scattered light issue is worse on the red side than it had been, although blue is a little better. NICFPS is still not available. There was a minor issue with Agile last night that is now fixed. The day staff is working on a DIS cart issue. Regarding ARCSAT, the telescope is available for this week. Things are mostly complete with the hardware and software upgrade that we did prior to the shutdown. The obs-specs are not quite up to speed on recovery and troubleshooting with the new system; Bill will be working with them this week to get them more familiar with the changes. He is also working on the integration of the FGCU Shelyak spectrometer that we have on loan. This is progressing nicely, and we hope to get on sky with that in a few weeks.

Moire asked if the reminder emails will go out earlier since the staff response time is likely to be slower? Answer: that is not currently the plan, although we could consider this if the users don't respond with enough lead time. Muk asked for a quantification of how much worse DIS red is? Russet replied that the halo of scattered light has shifted and gotten closer to core such that the boundary between the halo and the line profile core is indistinguishable. It looks like the spectrograph is out of focus or maybe has somewhat lower resolution. Blue has been in this state since the fall, but it has continued to get more compact such that the red end of the blue spectrum might be useable. The red side peak is ~ 50-60% of where it was. The blue peak is ~ 60% of where it was immediately after pumping on it last September and it is more than twice as good as it was in March. This degradation will have more of an effect on fainter targets. If one is looking at emission features one can still work with both sides of spectrograph. For absorption features it will be harder. We're not sure if the effect is linear so it could be that users will need to expose roughly twice as long in the red. Users are asked to provide feedback concerning the impact of the DIS degradation on their observations.


APO and the COVID-19 virus

  • Return to science. This was covered above; science operations resume tonight. Users are asked to a) be patient with the observing specialists, and b) reply to the reminder emails in a timely fashion.
  • Impact on summer shutdown. We will not be realuminizing the primary mirror at Kitt Peak this summer as originally scheduled, so our summer shutdown will be shorter. A group is meeting directly after this call to decide on the dates. Ben will contact the institutional schedulers if there is additional time allocate to the institutions once the new shutdown schedule is determined. Note added in proof: the summer shutdown dates will be July 6-14, with a scheduled return to science on July 15.
  • Impact on class trips. The site is still operating with the restrictions of no visitors to the site, so class training trips that were scheduled for June and July will be conducted remotely. It is unclear how long the visitors ban will be in place; anyone planning to travel to the site this fall should discuss this with the site staff ASAP.

Q2 3.5m scheduling:

All OPEN/DD time that had been remaining in Q2 has been allocated. Hopefully people who lost observing time due to the COVID-19 shutdown were able to recoup some time. We appreciate the numerous requests that were received and we were able to honor most of them, albeit not always the full request.


Q2 ARCSAT scheduling:

The Q2 schedule is full except for a few nights this week.


Q3 3.5m scheduling:

We have received all institutional requests so we have everything we need to get started on building the schedule. With the shorter summer shutdown we are likely to have more available dark time than we anticipated.


KOSMOS Update:

UW is working on a return to in-person research. They are currently in a "modified Phase 1" but can’t get the infection level low enough to change it. In the mean time the TEG is doing prep work to get it going once they are allowed to return to the lab. They have some reflective slits but to get more requires the reopening of the UW nanofab facility. To fully address the dewar problem requires cross-instrument calibration and coordination with multiple labs, which cannot happen yet.


ARCTIC Diffuser Rotation

The Preliminary Design Review was held a few weeks ago and several outstanding issues have been addressed. They are now getting quotes and ordering parts.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: There is no update this month. Nancy will request one for next month.
  • Getting narrow-band filters for ARCTIC: Nancy has finalized quotes from vendor; will place order.
  • Science Communications Interns: The two interns have been working on the following over the past several months: science highlights (posted to social media), staff highlights (posted to social media), new brochure design, and reviewing the web site. One idea they wanted to pitch to this group is that of an online hangout or lightning talks of science results that could be posted on our YouTube channel (and linked to from our web site once the new one gets deployed). This idea was well received and it was suggested that they look at the NRAO/Ask an Astronomer series on Facebook and Twitter and/or try to model this as a podcast type of discussion. It was also suggested that maybe the interns could start with a few interviews of NMSU people as examples for other institutions.

ACTION ITEMS:

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN.

New action items from this meeting:

  • UC reps: ask users for feedback on DIS performance.
  • UC reps: ask users to be patient with the night staff, and to respond to the reminder emails in a timely fashion.

All Other Business

None.


Next meeting:

If we continue alternating Mondays and Tuesdays each month to avoid standing conflicts, a possible schedule is: Tue Jul 7, Mon Aug 3, Tue Sep 1, Mon Oct 5, Tue Nov 3, Mon Dec 7

Having heard no objections to this, Nancy will send out the corresponding Zoom meeting invitations. The next meeting will be on Tuesday July 7 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.